clarisa:

find a number whose fourth part exceeds 7 by as much as its fifth part is less than 11

jhonyy9:

let be this number x so then i ve understood it correct \[x*\left( \frac{ 1 }{ 4 } \right) = 7 +y \] and \[x*\left( \frac{ 1 }{ 5 } \right) = 11 - y\] so and now just you need solve this system of equations hope helped